How can write energy of Hydrogen atom.
1
Expert's answer
2020-09-28T08:09:58-0400

There are a lot of ways how we can write energy of hydrogen atom. It depends on situation.

In molecular physics (macroscales) we don't have atoms of hydrogen, because hydrogen bounds with another hydrogen atom, and they make 2-atomic molecule. The energy of one molecule is

"\displaystyle E = \frac{5}{2}kT"

In atomic physics (microscales) energy of hydrogen is sometimes mixed up with the energy of electron in it. In this case, binding energy of electron in ground state(that is a sum of electron's potential and kinetic energies) can be written according to Bohr's model:

"\displaystyle E_1 = -\frac{me^4}{2 \hbar^2} = -13.6 \; eV."

We can also consider the rest mass of hydrogen. The full energy in this case is

"E = m_ec^2 + m_pc^2 = (m_e+m_p)c^2"

However, there are very few situation where this formula can be used, because this is energy released after total distraction (annihilation) of atom.